# LLM DNA: Tracing Model Evolution via Functional Representations

### v1

@@ 5.3 Overall Pipeline
This section presents the end-to-end DNA extraction pipeline based on the training-free algorithm. The procedure is detailed in Algorithm 1 and visualized in Figure 1. We begin by selecting a sample input set $\mathcal{S}_{t}$ of $t$ prompts drawn from real-world datasets. Each input $x_{i}$ is fed to an LLM $f$ to generate a textual response $y_{i}$ (line 4), which is then encoded into a fixed-size semantic vector $e_{i}$ using a sentence-embedding model (line 5). These $t$ embedding vectors are then concatenated to form a single, high-dimensional functional representation $E_{f}$ for the LLM (line 6). Finally, a pre-computed random Gaussian projection $A$ (line 1) linearly projects this high-dimensional representation $E_{f}$ into a low-dimensional vector $\tau_{f}\in\mathbb{R}^{L}$ (line 7), referred to as the DNA of the LLM.
@@
Algorithm 1 LLM DNA Extraction Pipeline
@@ 6 Experiments
This section presents the experimental details for constructing LLM DNAs and validates their effectiveness across multiple applications. Specifically, Section 6.1 presents datasets and models, with model details, hyperparameter setting, and environment included in Appendix B; Section 6.2 verifies that DNAs capture relationships among LLMs at scale; Section 6.3 evaluates DNAs for model routing; and Section 6.4 analyzes the stability of DNAs across heterogeneous datasets via a Mantel test. Furthermore, in Section 6.5 and Appendix C.1 (detailed version), we construct an evolutionary tree of LLMs purely based on DNAs that approximately matches the real-time evolution of LLMs.
@@ 6.1 Experimental Setting
Datasets. We evaluate across a diverse set of datasets covering question answering, commonsense reasoning, and general natural language understanding. Specifically, the pipeline integrates six widely used benchmarks, including SQuAD (Rajpurkar et al., 2016), CommonsenseQA (Talmor et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2020), ARC-Challenge (Clark et al., 2018), and MMLU (Hendrycks et al., 2021). DNA extraction employs a mixture of all six datasets, with each dataset contributing 100 random samples. Besides, EmbedLLM (Zhuang et al., 2024) dataset is also used to evaluate model routing, though not involved in DNA extraction. For each dataset, prompts are drawn from their designated text fields (e.g., question, sentence, or ctx).
LLMs. This study spans 305 LLMs released by 153 organizations (e.g. Qwen, Llama, etc.) on Hugging Face. The evaluation focuses on text-generative models, covering both decoder-only and encoder-decoder architectures, with parameter scales ranging from 100M to 70B. Both instruction-tuned or base models are included in the evaluation. Text embeddings are extracted with the state-of-the-art open-source encoder Qwen/Qwen3-Embedding-8B. More model details are presented in Appendix B.
@@ 6.2 Detect LLM Relationships
We first verify that DNA aligns with the findings of Zhu et al. (2025) on a limited set of LLMs. We then scale to 305 LLMs, showing that DNA accurately detects inter-model relationships. Finally, we visualize the DNA distribution for all 305 LLMs and highlight several related but under-documented model pairs.
@@
Figure 3: Mantel test between DNA extracted from two disjoint datasets. Each point represents a single pair of models, plotted by their distance in the first dataset versus their distance in the second, showing a strong correlation ($\text{Pearson-R}=0.7797$) and high statistical significance ($\text{P-value}=0.0000$).
Table 1 extends this experiment to 305 models. Using the official Hugging Face relationship (the “Model Tree”) as ground truth for correlated models, we obtain 83 correlated pairs. We additionally sample 83 random pairs and treat them as independent. The resulting dataset is split 8:2 into training and test sets. Because no existing method operates reliably across such heterogeneous models with distinct structures, we implement two baselines alongside our approach: Random—uniformly guessing independent or correlated; Greedy—predicting that all within-organization pairs are correlated and cross-organization pairs are independent; and DNA—an SVM with an RBF kernel trained on LLM DNAs. From Table 1, two observations emerge. First, DNA significantly outperforms the baselines and achieves a high AUC of $0.957$, indicating accurate detection of relationships among heterogeneous LLMs. Second, DNA attains higher recall than precision; this does not necessarily indicate a flaw in DNA, but rather suggests the presence of undocumented relationships in the official Model Tree. A detailed case study of the misclassified pairs appears in Appendix C.2.
@@
Method Precision Recall F1-Score AUC
Random 0.508 ±0.041 0.529 ±0.047 0.518 ±0.043 0.508 ±0.044
Greedy 0.764 ±0.083 0.349 ±0.000 0.478 ±0.016 0.618 ±0.024
DNA 0.858 ±0.072 0.916 ±0.102 0.882 ±0.066 0.957 ±0.044
@@ 6.3 Model Routing
This section investigates how DNA can benefit model routing. We use the same dataset, data split, baselines, and metric as EmbedLLM (Zhuang et al., 2024). EmbedLLM learns a compact LLM representation during routing training. To compare the quality of DNA with EmbedLLM-learned embeddings, within the matrix factorization framework of EmbedLLM, we replace the LLM embedding layer and its subsequent linear layer with frozen DNAs. These DNAs are directly compared with the query embeddings to make routing decisions. Notably, 11 models reported by Zhuang et al. (2024) either no longer exist or failed to run; their results are therefore omitted. Hence, our numbers do not perfectly align with those in Zhuang et al. (2024).
Table 2 reports routing accuracy, defined by Zhuang et al. (2024) as the fraction of queries routed to an LLM that answers correctly. The results show that frozen DNA outperforms the embeddings learned by EmbedLLM. Although the improvement is not significant, DNA is extracted without using the EmbedLLM dataset, whereas EmbedLLM learns its embeddings on this task. This suggests substantial potential for applying DNA to more broad applications.
@@ 6.4 Stability of DNA on Heterogeneous Datasets
This section examines how the choice of sampled distribution $S_{t}$ influences the quality of the generated DNAs. We assess this effect with a Mantel test between DNAs extracted from two disjoint datasets. Specifically, we construct one dataset as the union of 100 samples from SQuAD and 100 samples from CommonsenseQA (squad_cqa) and another as the union of 100 samples from HellaSwag and 100 samples from Winogrande (hs_wg). We compute DNAs from the same set of models, then calculate the distances between randomly selected pairs of models, visualize them in Figure 3, and compute the Pearson correlation and its $p$-value. From Figure 3, we observe a strong correlation ($\text{Pearson }r>0.75$) between the two distance structures with very high statistical significance ($p<0.001$). These results indicate that the structure of DNAs does not depend on the sampled data from which it is extracted, indicating the stability of our DNA extraction approach.
@@ Appendix A Proof
A.1 Proof of Lemma 3.1
[…]
Proof.
A.2 Proof of Theorem 3.2
[…]
@@
A.3 Proof of Theorem 3.2
[…]
@@
A.4 Proof of Theorem 4
@@
We invoke the JL Lemma with this $\epsilon$. The lemma applies because, as shown in Lemma 3.1, $(\mathcal{F},d_{H})$ is a Hilbert space, even when $d_{H}$ is defined with the weighting function $\mu(x)$. The lemma guarantees the existence of a linear map $E:\mathcal{F}\to\mathbb{R}^{L}$, with dimension $L=O(\epsilon^{-2}\log K)$, that embeds the vector representations of the functions. For any finite set $\{f_{1},\dots,f_{K}\}\subset\mathcal{F}$, the following holds for any pair $f_{i},f_{j}$:
@@ B.1 Hyperparameters
Unless otherwise specified, we adopt the following default configuration. The DNA dimensionality is reduced to 128 using Gaussian Random Projection. Text embeddings are extracted with the state-of-the-art open-source encoder Qwen/Qwen3-Embedding-8B, using a maximum input length of 1024 tokens with cls pooling. This model is also used to encode questions in routing experiment for semantic consistency. For text generation, we set max_length = 1024, temperature = 0.7, top_p = 0.9, if these options are applicable. Models with fewer than 7B parameters are run in BF16/FP16 precision without quantization, while larger models are automatically quantized to 8-bit for memory efficiency. For instruction-tuned models, the provided chat template is applied and only the responses are used for DNA extraction. All other parameters follow HuggingFace defaults. All applicable experiments are conducted on five seeds with mean and standard derivation reported.
B.2 Enviroment

### v2

@@ 4.3 Overall Pipeline
This section presents RepTrace algorithm, a training-free end-to-end DNA extraction pipeline. The procedure is detailed in Algorithm 1 and visualized in Figure 1. We begin by selecting a sample input set $\mathcal{S}_{t}$ of $t$ prompts drawn from real-world datasets. Each input $x_{i}$ is fed to an LLM $f$ to generate a textual response $y_{i}$ (line 4), which is then encoded into a fixed-size semantic vector $e_{i}$ using a sentence-embedding model (line 5). These $t$ embedding vectors are then concatenated to form a single, high-dimensional functional representation $E_{f}$ for the LLM (line 6). Finally, a pre-computed random Gaussian projection $A$ (line 1) linearly projects this high-dimensional representation $E_{f}$ into a low-dimensional vector $\tau_{f}\in\mathbb{R}^{L}$ (line 7), referred to as the DNA of the LLM.
@@
Algorithm 1 RepTrace: LLM DNA Extraction Pipeline
@@ 5 Experiments
This section provides a rigorous empirical validation of LLM DNA as a robust functional fingerprint. Following our experimental setup across 305 models (§5.1, §B), we demonstrate DNA’s efficacy in relationship detection (§5.2) and training-free model routing (§5.3). We verify the representation’s stability through stress tests involving data heterogeneity (§5.4), synthetic random inputs (§5.5), chat template variations (§5.6), and fine-tuning intensity (§5.7). Systematic ablations characterize the influence of bottleneck embedding models (§C.3), pre-aggregation dimensions ppC.4), and DNA length LLC.5). To address technical edge cases, we analyze parameter-size bias (§C.6), same-tokenizer performance (§C.7), and relationship ”false positives” (§C.2). We conclude by leveraging these functional distances to reconstruct a phylogenetic tree of LLMs (§5.8, §C.1), providing a data-driven map of the field’s architectural evolution.
@@ 5.1 Experimental Setting
Datasets. We evaluate across a diverse set of datasets covering question answering, commonsense reasoning, and general natural language understanding. Specifically, the pipeline integrates six widely used benchmarks, including SQuAD (Rajpurkar et al., 2016), CommonsenseQA (Talmor et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2020), ARC-Challenge (Clark et al., 2018), and MMLU (Hendrycks et al., 2021). DNA extraction employs a mixture of all six datasets, with each dataset contributing 100 random samples. Besides, EmbedLLM (Zhuang et al., 2025) dataset is also used to evaluate model routing, though not involved in DNA extraction. For each dataset, prompts are drawn from their designated text fields (e.g., question, sentence, or ctx).
LLMs. This study spans 305 LLMs released by 153 organizations (e.g. Qwen, Llama, etc.) on Hugging Face. The evaluation focuses on text-generative models, covering both decoder-only and encoder-decoder architectures, with parameter scales ranging from 100M to 70B. Both instruction-tuned or base models are included in the evaluation. Text embeddings are extracted with the state-of-the-art open-source encoder Qwen/Qwen3-Embedding-8B by default, with performance comparison to sentence-transformers/all-mpnet-base-v2 and BAAI/bge-large-en-v1.5 presented. More model details are included in Appendix B.
@@ 5.2 Detect LLM Relationships
In this subsection, we validate DNA against Zhu et al. (2025), scale to 305 LLMs to demonstrate accurate relationship detection, and visualize the distribution to uncover undocumented model pairs.
@@
Figure 3: Mantel test between DNA extracted from two disjoint datasets. Each point represents a single pair of models, plotted by their distance in the first dataset versus their distance in the second, showing a strong correlation ($\text{Pearson-$r$}=0.7797$) and high statistical significance ($\text{$p$-value}<0.0001$).
Table 1 extends this experiment to 305 models. Using the official Hugging Face relationship (the “Model Tree”) as ground truth for correlated models, we obtain 83 correlated pairs. We additionally sample 83 random pairs and treat them as independent. The resulting dataset is split 8:2 into training and test sets. Three baselines are compared alongside our approach: Random—uniformly guessing independent or correlated; Greedy—predicting that all within-organization pairs are correlated and cross-organization pairs are independent; and PhyloLM (Yax et al., 2025)—genetic (Nei) distances to other LLMs in the training set are regarded as a signature vector. For PhyloLM and DNA, an SVM with RBF kernel is used for prediction. From Table 1, three observations emerge. First, DNA significantly outperforms the baselines and achieves a high AUC of $0.992$, indicating accurate detection of relationships among heterogeneous LLMs. Second, DNA attains higher recall than precision; this does not necessarily indicate a flaw in DNA, but rather suggests the presence of undocumented relationships in the official model tree. A detailed case study of the misclassified pairs appears in Appendix C.2. Third, DNA quality is robust to the choice of sentence-embedding model. Using BGE (0.3B) and MPNet (0.1B), despite their much smaller parameter sizes, yields comparable—and in this relation prediction task, even better—performance.
@@
Method Accuracy Precision Recall F1 AUC
Random 0.493 ±0.036 0.515 ±0.034 0.516 ±0.042 0.516 ±0.036 0.492 ±0.036
Greedy 0.593 ±0.018 0.759 ±0.081 0.329 ±0.000 0.458 ±0.014 0.606 ±0.023
PhyloLM + SVM 0.742 ±0.058 0.741 ±0.073 0.812 ±0.123 0.766 ±0.057 0.788 ±0.060
DNA (Qwen3) + SVM 0.919 ±0.048 0.898 ±0.049 0.957 ±0.072 0.925 ±0.047 0.979 ±0.027
DNA (BGE) + SVM 0.934 ±0.053 0.905 ±0.068 0.982 ±0.059 0.940 ±0.050 0.992 ±0.023
DNA (MPNet) + SVM 0.932 ±0.048 0.914 ±0.065 0.966 ±0.069 0.937 ±0.047 0.990 ±0.020
@@ 5.3 Model Routing
This section investigates how DNA can benefit model routing. We use the same dataset, data split, baselines, and metric as EmbedLLM (Zhuang et al., 2025). EmbedLLM learns a compact LLM representation during routing training. To compare the quality of DNA with EmbedLLM-learned embeddings, within the matrix factorization framework of EmbedLLM, we replace the LLM embedding layer and its subsequent linear layer with frozen DNAs. These DNAs are directly compared with the query embeddings to make routing decisions. Notably, 11 models reported by Zhuang et al. (2025) either no longer exist or failed to run; their results are therefore omitted. Hence, our numbers do not perfectly align with those in Zhuang et al. (2025).
Table 2 reports routing accuracy, defined by Zhuang et al. (2025) as the fraction of queries routed to an LLM that answers correctly. The results show that frozen DNA outperforms the embeddings learned by EmbedLLM. This result is notable because EmbedLLM explicitly learns representations on this routing dataset, whereas LLM DNA is training-free and task-agnostic. Achieving parity without task-specific optimization validates the quality of DNA and highlights its superior scalability: unlike EmbedLLM, our model supports arbitrary new models without retraining.
@@ 5.4 Stability of DNA on Heterogeneous Datasets
This subsection examines how the choice of sampled distribution $S_{t}$ influences the quality of the generated DNAs. We assess this effect with a Mantel test between DNAs extracted from two disjoint datasets. Specifically, we construct one dataset as the union of 100 samples from SQuAD and 100 samples from CommonsenseQA (squad_cqa) and another as the union of 100 samples from HellaSwag and 100 samples from Winogrande (hs_wg). We compute DNAs from the same set of models, then calculate the distances between randomly selected pairs of models, visualize them in Figure 3, and compute the Pearson correlation and its $p$-value. From Figure 3, we observe a strong correlation ($\text{Pearson }r>0.75$) between the two distance structures with very high statistical significance ($p<0.001$). These results indicate that the structure of DNAs does not depend on the sampled data from which it is extracted, indicating the stability of our DNA extraction approach.
5.5 DNA Extracted from Random Data
[…]
The results are reported in Table 3. Interestingly, the random-data setting even slightly outperforms the standard benchmark. This highlights that the DNA extraction method is largely insensitive to input-dataset bias. The key intuition is that DNA analyses (e.g., relationship prediction) target relative relationships among LLMs rather than absolute accuracy. Different datasets act as different “views” of the same model set: a suboptimal view may obscure some relations, but it is unlikely to fabricate spurious ones—for instance, it will not make two genuinely similar models appear unrelated. This observation also aligns with the stability we see under random projection.
Table 3: DNA (Qwen3) relation prediction performance (DNA extracted from random inputs)
Method Accuracy Precision Recall F1 AUC
PhyloLM + SVM 0.742 ±0.058 0.741 ±0.073 0.812 ±0.123 0.766 ±0.057 0.788 ±0.060
DNA + SVM 0.919 ±0.048 0.898 ±0.049 0.957 ±0.072 0.925 ±0.047 0.979 ±0.027
DNA (Rand) + SVM 0.949 ±0.048 0.931 ±0.053 0.977 ±0.062 0.952 ±0.048 0.987 ±0.038
5.6 Ablation Study for Chat Template
We study how chat templates influence DNA extraction for instruction-tuned models. Chat templates inject the special tokens and formatting used during instruction tuning to elicit conversational behavior. For an instruction-tuned model (e.g., Llama3-Instruct), we compare two protocols: (i) bypass the template and feed the raw prompt directly (completion-style), and (ii) apply the template and compute DNA using only the extracted assistant’s response. For non-chat (base/completion) models, the extraction pipeline is unchanged. We report relationship prediction performance under the same setting as Table 4. Removing the chat template consistently improves accuracy by 1–3% across encoders, suggesting that for cross-family comparisons involving both chat-tuned and completion models, omitting templates yields a fairer and more direct functional comparison by evaluating chat-tuned models through the same completion-style interface as their base counterparts.
Table 4: DNA relation prediction performance: with vs without chat templates
Encoder Template Accuracy Precision Recall F1 AUC
Qwen3 w/ template 0.919 ±0.048 0.898 ±0.049 0.957 ±0.072 0.925 ±0.047 0.979 ±0.027
w/o template 0.930 ±0.045 0.916 ±0.067 0.961 ±0.071 0.935 ±0.044 0.994 ±0.021
MPNet w/ template 0.932 ±0.048 0.914 ±0.065 0.966 ±0.069 0.937 ±0.047 0.990 ±0.020
w/o template 0.958 ±0.047 0.943 ±0.060 0.982 ±0.059 0.960 ±0.047 0.994 ±0.016
BGE w/ template 0.934 ±0.053 0.905 ±0.068 0.982 ±0.059 0.940 ±0.050 0.992 ±0.023
w/o template 0.945 ±0.054 0.929 ±0.073 0.974 ±0.066 0.948 ±0.052 0.992 ±0.023
5.7 Effect of Fine-tuning on the DNA Values
This subsection quantifies how fine-tuning changes DNA values. Starting from Llama3-8B-Instruct, we fully fine-tune all parameters on subsets of the NVIDIA OpenMathInstruct-2 dataset (Toshniwal et al., 2024) with sizes 10,100,1000,10000{10,100,1000,10000}. Since Llama3 was released before OpenMathInstruct-2, this dataset appears in neither its pretraining corpus nor our DNA extraction set. Figure 5 compares the DNAs of the original and fine-tuned models and reports their $L_{2}$ distances to the original DNA.
Figure 5: Shift of DNA values when (full-parameter) fine-tuning Llama3-8B-Instruct with OpenMathInstruct-2 subsets of size $n$
We draw two conclusions. First, the DNA distance increases with the amount of fine-tuned data, showing a monotonic but nonlinear trend. Specifically, the distance rises sharply after fine-tuning on a tiny external subset (e.g., the first 10 samples causes $L_{2}=0.69$) and then grows more gradually (up to 0.80 with 10,000 samples) as additional samples are learned. Second, the global DNA structure remains stable: most subsequences preserve their dark/light patterns with only minor variation, suggesting that fine-tuning alters local traits while leaving inherited, intrinsic behaviors largely intact.
@@ 6 Conclusion
Reproducibility Statement
To facilitate reproducibility, we release our source code at https://github.com/Xtra-Computing/LLM-DNA and an interactive demo at https://dna.xtra.science. We also provide a PyPI package, available at https://pypi.org/project/llm-dna, which can be installed via pip install llm-dna. Regarding theoretical and experimental details, proofs appear in Appendix A, data specifications in Section 5.1, and hyperparameters, model details, and licenses in Appendix B.
@@
Acknowledgement
This research/project is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority.
@@ Appendix A Proof
A.1 LLM Properties: Inheritance and Genetic Determinism
Before discussing the properties, we begin by defining the evolution of an LLM.
Definition A.1 (Evolution).
Let $\mathcal{F}$ be the LLM function space in Definition 3.1, and let $d_{H}$ be the Hilbert distance on $\mathcal{F}$. For a threshold $\delta_{H}>0$, an evolution is a function $\mathcal{E}:\mathcal{F}\rightarrow\mathcal{F}$ such that for any $f\in\mathcal{F}$,
$d_{H}\big(f,\mathcal{E}(f)\big)<\delta_{H}.$
If $f_{2}=\mathcal{E}(f_{1})$, we say $f_{2}$ is evolved from $f_{1}$ within $\delta_{H}$.
This notion covers a wide range of operations that induce a bounded shift in an LLM’s functional behavior, including but not limited to fine-tuning, distillation, and reinforcement learning, as they modify the LLM within a bounded Hilbert-distance threshold.
Theorem A.2 (Inheritance).
Let an LLM $f_{2}\in\mathcal{F}$ be derived from an LLM $f_{1}\in\mathcal{F}$, with corresponding DNAs $\tau_{f_{2}}$ and $\tau_{f_{1}}$. For any desired DNA proximity $\epsilon_{\tau}>0$, there exists a Hilbert distance threshold $\delta_{H}>0$ such that if $d_{H}(f_{1},f_{2})<\delta_{H}$, then $d_{\tau}(\tau_{f_{1}},\tau_{f_{2}})<\epsilon_{\tau}$.
@@ A.1 LLM Properties: Inheritance and Genetic Determinism
Theorem A.3 (Genetic Determinism).
Let $f_{1},f_{2}\in\mathcal{F}$ be two LLMs with corresponding DNAs $\tau_{f_{1}},\tau_{f_{2}}\in\mathcal{D}$. For any desired functional proximity $\epsilon_{H}>0$, there exists a DNA distance threshold $\delta_{\tau}>0$ such that if $d_{\tau}(\tau_{f_{1}},\tau_{f_{2}})<\delta_{\tau}$, then $d_{H}(f_{1},f_{2})<\epsilon_{H}$.
@@
A.2 LLM Functional Space is a Hilbert Space
Lemma A.4 (LLM Functional Space is a Hilbert Space).
Let $\mathcal{F}$ be the space of LLM functions mapping from the finite input space $\mathcal{S}_{m}=\{x_{1},\dots,x_{N}\}$ to the vector space of outputs $\mathcal{O}:=\mathbb{R}^{N}$. Let $\mu:\mathcal{S}_{m}\to[0,1]$ be a probability distribution on the input space, assigning a probability $\mu(x_{i})$ to each sequence such that $\sum_{i=1}^{N}\mu(x_{i})=1$. The pair $(\mathcal{F},d_{H})$ forms a Hilbert space, where the distance metric $d_{H}$ is defined as: $d_{H}(f_{1},f_{2})=\sqrt{\sum_{i=1}^{N}\mu(x_{i})\|f_{1}(x_{i})-f_{2}(x_{i})\|_{2}^{2}}$
Proof.
A.3 Proof of Theorem 3.2
@@ A.2 LLM Functional Space is a Hilbert Space
We invoke the JL Lemma with this $\epsilon$. The lemma applies because, as shown in Lemma A.4, $(\mathcal{F},d_{H})$ is a Hilbert space, even when $d_{H}$ is defined with the weighting function $\mu(x)$. The lemma guarantees the existence of a linear map $E:\mathcal{F}\to\mathbb{R}^{L}$, with dimension $L=O(\epsilon^{-2}\log K)$, that embeds the vector representations of the functions. For any finite set $\{f_{1},\dots,f_{K}\}\subset\mathcal{F}$, the following holds for any pair $f_{i},f_{j}$:
@@
A.5 Proof of Corollary 3.2
[…]
Proof.
Since Lemma A.4 shows that $\mathcal{F}$ is a Hilbert space, the JL lemma applies. Invoking a standard formulation of the lemma (e.g., Theorem 2.1 in Dasgupta & Gupta (2003)) for the set of vector representations of LLMs in $\mathcal{F}_{K}$ yields the desired projection and associated probability bound. ∎
@@ B.1 Hyperparameters
Unless otherwise specified, we adopt the following default configuration. The DNA dimensionality is reduced to 128 using Gaussian Random Projection by default. Text embeddings are extracted with the state-of-the-art open-source encoder Qwen/Qwen3-Embedding-8B, using a maximum input length of 1024 tokens with cls pooling. This model is also used to encode questions in routing experiment for semantic consistency. For text generation, we set max_length = 1024, temperature = 0.7, top_p = 0.9, if these options are applicable. Models with fewer than 7B parameters are run in BF16/FP16 precision without quantization, while larger models are automatically quantized to 8-bit for memory efficiency. For instruction-tuned models, the provided chat template is applied and only the responses are used for DNA extraction. All other parameters follow HuggingFace defaults. All applicable experiments are conducted on five seeds with mean and standard deviation reported.
B.2 Environment